OpenStudy (anonymous):

How do you solve -5/4k+2/9=-6+3/8k?

OpenStudy (anonymous):

\[\frac{-5}4 k + \frac{2}9 = -6 + \frac{3}8k\]?

OpenStudy (anonymous):

yesss?

OpenStudy (anonymous):

make the denominators common \[-\frac{90}{72}k+\frac{16}{72} = -\frac{432}{72} + \frac{27}{72}k\]can you go on from here?

OpenStudy (anonymous):

yes i think so,, thank you